OpenStudy (i_love_my_nieces):

y = (45,000) (.08)^15

OpenStudy (i_love_my_nieces):

Simplify

OpenStudy (c1ndrella78):

y=45000 (.08)^15 First simplify (.08)^15 What would that =?

OpenStudy (i_love_my_nieces):

I got 0

OpenStudy (i_love_my_nieces):

These are my choices

OpenStudy (denonakavro):

just use a calculator 45000x 0.8^15 45000x 0.035184 =1583.3

OpenStudy (i_love_my_nieces):

So a?

OpenStudy (denonakavro):

Are you sure your question is correct ?

OpenStudy (mathstudent55):

The decay starts at 100%. After 1 year, it loses 2%, so it is now 98%. As it goes on, each year, you have 98% of the previous year. The equation is: y = 45000(0.98)^15

OpenStudy (c1ndrella78):

You should be able to just add it in your calculator, Check these out. http://www.regentsprep.org/Regents/math/ALGEBRA/AE7/ExpDecayL.htm http://www.regentsprep.org/Regents/math/ALGEBRA/AE7/ExpDecayL.htm You got the right equation

OpenStudy (i_love_my_nieces):

This is what I got 33235.609619

OpenStudy (mathstudent55):

The population starts at 100%. After 1 year, it loses 2%, so it is now 98% of the population. As it goes on, each year, you have 98% of the previous year. The equation is: y = 45000(0.98)^15

OpenStudy (mathstudent55):

Correct.

OpenStudy (i_love_my_nieces):

So b

OpenStudy (mathstudent55):

Yes.